# Bridging Developer Instructions and Code Completion Through Instruction-Aware Fill-in-the-Middle Paradigm

### v1

@@ Abstract.
Large Language Models (LLMs) have significantly advanced code completion, yet they often fail when the developer’s intent is underspecified in the code context. In practice, developers frequently add natural language instructions (e.g., comments) to clarify their intent. However, exisintg code LLMs applied for code completion systems merely undergoes a fill-in-the-middle (FIM) pre-training, which struggle to leverage this information effectively due to the lack of instruction-like training data. Existing instruction-tuning techniques, which improve instruction-following in general code generation, paradoxically degrade FIM performance, forcing a trade-off between instruction-following and infilling capabilities. To address this gap, we introduce Instruction-aware Fill-in-the-Middle (IFIM), an instruction-tuning method specifically designed to enhance FIM code completion models. IFIM extends the conventional FIM training objective by incorporating an explicit instruction section into the input, enabling the model to learn from (prefix, instruction, suffix) triplets. This approach allows the model to effectively leverage developer-provided directives while preserving its core completion abilities when no instructions are present. To facilitate this, we constructed a large-scale dataset by using GPT-4o to generate concise, intent-focused instructions for code infilling examples. We evaluated IFIM by applying it to two popular base models, Deepseek-Coder and Qwen2.5-Coder, on the benchmarks derived from HumanEval-infilling and RepoMasterEval. The results demonstrate that IFIM significantly improves instruction-following capabilities, boosting the Pass@1 score from 84.6% to 93.6% on HumanEval-infilling. Moreover, this enhancement does not compromise the models’ original performance on FIM code completion tasks with no instructions provided.
copyright: noneconference: 48th International Conference on Software Engineering; April 12–18, 2026; Rio De janeiro, Brazil
@@ 1. Introduction
Modern Integrated Development Environments (IDEs) rely on code completion to accelerate development (Wang et al., 2023): given a partial program, a model predicts the missing sections and displays them to the programmer. Large Language Models (LLMs) have pushed this feature from keyword-level suggestions to full-function generation, which adapts to different coding styles and languages with remarkable accuracy. As a result, coding assistants, with LLM-powered code completion as a key feature, have become a buzzword in the developer community, such as Cursor, Github Copilot, and Trea.
@@
We evaluate IFIM on two code completion benchmarks IHumaneval and IRepoMasterEval, derived respectively from Humaneval-infilling (Fried et al., 2022) and RepoMasterEval (Wu et al., 2024) with code completion instructions added to their data samples. To investigate the impact of IFIM training on widely used code completion LLMs, we experiment with two base code LLMs: Deepseek-Coder (DeepSeek-AI et al., 2024) and Qwen2.5-Coder (Hui et al., 2024). Our comprehensive experiments demonstrate that IFIM achieves remarkable improvements in instruction-following capability while maintaining the LLMs’ ability to complete the code without instructions. Specifically, given the same instructions, our IFIM-trained variants can achieve better performance than the base model, with Pass@1 score averagely improved from 87.8% to 94.7% on IHumaneval and from 14.7% to 20.7% on IRepoMasterEval. Crucially, when no instructions are provided, IFIM models either match or slightly exceed their base counterparts’ performance. For example, on the IHumaneval benchmark, IFIM improves Deepseek-Coder from 68.6% to 78.2% and Qwen2.5-Coder from 76.0% to 76.3%. Moreover, we also conducted a series of ablation experiments to analyze the optimal position for inserting the instruction component, the ratio of IFIM processed data, and the advantages of using a standalone instruction component. These results position IFIM as a practical, backward-compatible upgrade path for existing LLM-based code completion systems.
@@ 2.1. Fill-in-the-Middle Code Completion
The Fill-in-the-Middle (FIM) objective (Bavarian et al., 2022) was proposed to extend the conventional left-to-right, causal language-modeling paradigm with a crucial infilling capacity during pre-training. It grants the model the ability to generate a contiguous span of text that is missing from the middle of a document while conditioning on both the preceding (prefix) and the succeeding (suffix) context. Code completion is the natural test-bed for this capability, as developers routinely need to insert new logic between already-written lines, such as the body of a newly declared function, missing import statements, or boilerplate loops. An illustrative example of this is shown in Figure 1. Because the content that follows the developer’s cursor is already known, a model that can exploit this suffix context is strictly more informative and powerful than one restricted to a left-only prefix.
@@ 2.2. Base Code LLMs are Sub-Optimal Instruction Followers
Base code LLMs are pre-trained on a vast data corpus, the vast majority of which consists of code files scraped from public repositories, such as GitHub. Consequently, during code completion, they fill the gap with whatever appears most likely based on patterns learned from the training data. This extends to natural language instructions written as comments; the model “follows” them by predicting the most probable code sequence found after similar comments in its training data. However, the developer’s true, fine-grained intent while coding is a signal that is largely absent from the public code repositories used for pre-training. Much of that intent never reaches the committed source: it may surface as a fleeting search-engine query, a short Slack thread with a teammate, or even remain unspoken in the developer’s mind. The comments that are finally documented in the code are, therefore, an insufficient proxy for instructions. Moreover, prior studies (Sun et al., 2022; Shi et al., 2022) have shown that the comments in source code are very noisy, with only a small fraction of comments actually describing the functionality of the code. The majority serve other purposes, such as licensing information, TODOs, or internal team coordination, making them poor indicators of intent. Relying on them alone leaves the model under-conditioned for real-world scenarios where intent is expressed ad-hoc. To close the gap, we need to supplement the model’s strong FIM foundation with targeted instruction alignment that empowers it to interpret explicit natural-language instructions.
@@ 2.3. Conversational LLMs are Sub-Optimal Code Completers
Conversational LLMs, such as GPT-4o, Claude-3, and DeepSeek-R1, have achieved remarkable performance on code-related dialog tasks. These include answering debugging questions, generating entire programs from natural-language specifications, and refactoring code snippets. In these scenarios, the model receives a textual instruction and produces a helpful, conversational response that mixes text and code. This capability is cultivated through an instruction tuning stage, where models learn to follow human instructions and dialog formats, often through supervised learning on instruction-response pairs (Wei et al., 2021). However, the datasets used for this purpose, such as Magicoder (Wei et al., 2023) and Wizardcoder (Luo et al., 2023), primarily emphasize generating complete, assistant-style outputs rather than context-aware code insertions. This focus comes at a cost: the training process largely overlooks the strict contextual and logical consistency required by FIM code completion, where new code must integrate seamlessly with its surrounding prefix and suffix. As a result, when applied to FIM tasks, these powerful conversational models exhibit various failure modes, such as repeating the suffix or generating irrelevant content (Ahmad et al., 2025; Gong et al., 2024; Pan et al., 2024). An FIM-trained base model like Codegemma-7b (Team et al., 2024) can easily outperform the much larger conversational LLM, GPT-4o, on real-world code completion benchmarks (Pan et al., 2024). Consequently, the LLMs deployed in production code completion systems are typically pre-trained base models, rather than their instruction-tuned counterparts.
@@ 3.3. Usage for Code Completion
An IFIM-trained LLM enables a developer to guide code completion by embedding instructions directly within their code. Since plain text would violate the programming language’s syntax, we propose writing instructions as specially marked inline comments. An example of this is shown in Figure 3. In this example, a developer using Python could write “#!filter out negative numbers” as the instruction. The ‘#!’ prefix acts as a unique marker that distinguishes an IFIM instruction from a regular comment. This allows client-side tools, like IDE plugins, to build user-friendly features around it. For example, an IDE could be configured to automatically remove the instructional comment once the generated code is accepted by the user. When a code completion request is made, the code completion system parses the code around the cursor to extract the prefix and suffix, and identifies the specially marked comment as the instruction. These three components, prefix, suffix, and instruction, are then formatted into the required IFIM input sequence. The IFIM-trained LLM receives this input and generates the code to fill the middle, which is then presented to the developer as a completion suggestion.
@@ 4. Experiment Setup
  • RQ4: Does representing instructions as a dedicated component outperform using instructions as inline comments?
@@ 4.2. Code LLMs
We evaluate our approach using two popular code LLMs as base models: Deepseek-Coder (Guo et al., 2024) and Qwen-2.5-Coder (Hui et al., 2024). Each model family includes a base version (pretrained with FIM objectives) and an instruction-tuned version (further fine-tuned on proprietary instruction datasets). In our experiments, we apply IFIM training to the base versions since they are the LLMs used for code completion.
@@ 4.3. Evaluation Benchmarks
Notably, these two benchmarks are not natively paired with code completion instructions. To evaluate the instruction-following capability of the models, we synthesize instructions for each sample using OpenAI GPT-4.1 (gpt-4.1-2025-04-14) with the same prompt employed in our instruction synthesis process. To further ensure the quality of the instructions, we manually revise those of low quality. Given the large sample sizes, we randomly select statistically representative subsets from each benchmark (95% confidence level, 5% margin of error, 50% population proportion), resulting in 312 and 256 samples for HumanEval-infilling and RepoMasterEval, respectively. The manual filtering process involves two authors independently reviewing all generated instructions and revising those that inadequately describe the desired code functionality. The authors achieve an inter-rater agreement of 0.72 and collaboratively resolve disagreements through discussion. The final curated instruction benchmarks are respectively named as IHumanEval and IRepoMasterEval (IRME), which are used in the experiments.
@@ 5.1. RQ1: Effectiveness of IFIM
In this RQ, we systematically evaluate the effectiveness of IFIM, where we aim to assess two aspects: (1) whether IFIM can improve the instruction-following ability of the LLMs, and (2) whether IFIM can maintain the code completion ability when no instruction is given. Specifically, for the two base LLMs, Deepseek-coder and Qwen2.5-coder, we conduct IFIM training and obtain corresponding IFIM variants. The obtained models are evaluated with the two benchmarks, IHumanEval and IRME, under two settings: (1) with instruction and (2) without instruction. In the foIRMEr setting, we pair each completion request with its corresponding instruction, where the instructions for base models are written as a comment appended to the end of the prefix, while the instructions for IFIM variants are constructed following the IFIM mode. In the latter setting, we use pure prefixes and suffixes without any modification. The Pass@1 score is computed to assess the quality of the completion.
@@
w/ ins. Deepseek-Coder base PMS 84.6% 10.9%
IFIM PIMS 93.6% 21.1%
Qwen2.5-Coder base PSM 91.0% 18.4%
IFIM PSIM 95.8% 20.3%
w/o ins. Deepseek-Coder base PMS 68.6% 7.4%
IFIM PIMS 78.2% 16.0%
Qwen2.5-Coder base PSM 76.0% 10.2%
IFIM PSIM 76.3% 13.3%
@@ 5.2. RQ2: Impact of IFIM Mode
Notably, the I-before-M modes (i.e., placing the Instruction component immediately before the Middle component) consistently achieve the highest instruction-following performance compared with other IFIM modes. For Deepseek-Coder, the PIMS mode is the top perfoIRMEr on both IHumanEval (93.6%) and IRME (21.1%). Similarly, for Qwen2.5-Coder, the PSIM mode achieves the best scores on both IHumanEval (95%) and IRME (20.3%). This demonstrates that placing the instruction immediately before the code to be generated most effectively enhances the LLMs’ understanding and execution of the given instructions. We therefore adopt I-before-M as the default mode for IFIM training.
@@ 5.4. RQ4: Impact of Standalone Instruction Component
In this RQ, we seek to investigate the impact of our newly introduced Instruction component, compared with simply using instruction as inline comments. Specifically, we derive two training datasets from our instruction dataset: one processed using the IFIM format and the other by appending the instruction as an inline comment to the prefix. Using these two derived datasets, we train each base model (Deepseek-Coder and Qwen2.5-Coder) and correspondingly obtain two variants. Finally, we have two base models, each with one IFIM variant and one experimental variant for inline comments (named as CFIM, short for Comment-trained FIM). Since the instruction is appended as a part of the prefix, the inference for CFIM variants follows the same mode as their base models. We evaluate the performance of these models on IHumanEval-infilling and RepoMasterEval. During evaluation, the IFIM models are provided with instructed inputs, and the base model and CFIM models are provided with commented inputs (i.e., instructions are appended to the prefix as an inline comment). We compute the Pass@1 scores to measure the correctness of the completion.
@@
The results presented in Table 4 reveal that simply adding instructions as inline comments is an ineffective and often detrimental strategy. For Deepseek-Coder, this approach results in degraded performance, with Pass@1 scores dropping from 84.6% to 79.0% on IHumanEval and from 10.9% to a mere 4.3% on IRME. The Qwen2.5-Coder CFIM variant showed a similar, though less pronounced, negative effect, with performance decreasing on IHumanEval and remaining stagnant on IRME.
This confirms a hypothesis: LLMs can be confused when instructional text is mixed with code as comments, negatively impacting their primary code completion capabilities. The stark contrast between the significant gains from IFIM and the performance degradation from CFIM suggests that the structural separation of instructions from the code, as implemented in IFIM, is essential for the model to properly disambiguate, understand, and benefit from the provided guidance.

### v2

@@ Abstract.
Large Language Models (LLMs) have significantly advanced code completion, yet they often fail when the developer’s intent is underspecified in the code context. To address this, developers usually add natural language instructions (e.g., comments) into the code context to clarify their intent. However, existing code LLMs applied for code completion systems merely undergo a fill-in-the-middle (FIM) pre-training, which struggles to leverage this information effectively due to the lack of instruction-like training data. Existing instruction-tuning techniques, which improve instruction-following in general code generation, paradoxically degrade FIM performance, forcing a trade-off between instruction-following and infilling capabilities. To address this gap, we introduce Instruction-aware Fill-in-the-Middle (IFIM), an instruction-tuning method specifically designed to enhance FIM code completion models. IFIM extends the conventional FIM training objective by incorporating an explicit instruction section into the input, enabling the model to learn from (prefix, instruction, suffix) triplets. This approach allows the model to effectively leverage developer-provided directives while preserving its core completion abilities when no instructions are present. To facilitate this, we constructed a large-scale dataset by using GPT-4o to generate concise, intent-focused instructions for code infilling examples. We evaluated IFIM by applying it to two popular base models, Deepseek-Coder and Qwen2.5-Coder, on the benchmarks derived from HumanEval-infilling and RepoMasterEval. The results demonstrate that IFIM significantly improves instruction-following capabilities, boosting the Pass@1 score from 84.6% to 93.6% on HumanEval-infilling. Moreover, this enhancement does not compromise the models’ original performance on FIM code completion tasks with no instructions provided.
copyright: none
@@ 1. Introduction
Modern Integrated Development Environments rely on code completion to accelerate development (Wang et al., 2023): given a partial program, a model predicts the missing sections and displays them to the programmer. Large Language Models (LLMs) have pushed this feature from keyword-level suggestions to full-function generation, which adapts to different coding styles and languages with remarkable accuracy. As a result, coding assistants, with LLM-powered code completion as a key feature, have become a buzzword in the developer community, such as Cursor, Github Copilot, and Trea.
@@
We evaluate IFIM on two code completion benchmarks IHumaneval and IRepoMasterEval, derived respectively from Humaneval-infilling (Fried et al., 2022) and RepoMasterEval (Wu et al., 2024) with human-annotated code completion instructions added to their data samples. To investigate the impact of IFIM training on widely used code completion LLMs, we experiment with two base code LLMs: Deepseek-Coder (DeepSeek-AI et al., 2024) and Qwen2.5-Coder (Hui et al., 2024). Our experiments demonstrate that IFIM achieves remarkable improvements in instruction-following capability while maintaining the LLMs’ ability to complete the code without instructions. Specifically, given the same instructions, our IFIM-trained variants can achieve better performance than the base model, with Pass@1 score averagely improved from 87.8% to 94.7% on IHumaneval and from 14.7% to 20.7% on IRepoMasterEval. Crucially, when no instructions are provided, IFIM models either match or slightly exceed their base counterparts’ performance. For example, on the IHumaneval benchmark, IFIM improves Deepseek-Coder from 68.6% to 78.2% and Qwen2.5-Coder from 76.0% to 76.3%. Moreover, we also conducted a series of ablation experiments to analyze the optimal position for inserting the instruction component, the ratio of IFIM processed data, the advantages of using a standalone instruction component and the alignment between instructions and generated completions. These results position IFIM as a practical, backward-compatible upgrade path for existing LLM-based code completion systems.
@@ 2.1. Fill-in-the-Middle Code Completion
The FIM objective (Bavarian et al., 2022) was proposed to extend the conventional left-to-right, causal language-modeling paradigm with a crucial infilling capacity during pre-training. It grants the model the ability to generate a contiguous span of text that is missing from the middle of a document while conditioning on both the preceding (prefix) and the succeeding (suffix) context. Crucially, this new infilling capability does not compromise the model’s original performance in standard left-to-right generation. Code completion is the natural test-bed for this capability, as developers routinely need to insert new logic between already-written lines, such as the body of a newly declared function, missing import statements, or boilerplate loops. An illustrative example of this is shown in Figure 1. Because the content that follows the developer’s cursor is already known, a model that can exploit this suffix context is strictly more informative and powerful than one restricted to a left-only prefix.
@@ 2.2. Base Code LLMs are Sub-Optimal Instruction Followers
Base code LLMs are pre-trained on a vast data corpus, the vast majority of which consists of code files scraped from public repositories, such as GitHub. Consequently, during code completion, they fill the gap with whatever appears most likely based on patterns learned from the training data. This extends to natural language instructions written as comments; the model “follows” them by predicting the most probable code sequence found after similar comments in its training data. However, the developer’s true, fine-grained intent while coding is a signal that is largely absent from the public code repositories used for pre-training. Much of that intent never reaches the committed source code: it may surface as a fleeting search-engine query, a short Slack thread with a teammate, or even remain unspoken in the developer’s mind. The comments that are finally documented in the code are, therefore, an insufficient proxy for instructions. Moreover, prior studies (Sun et al., 2022; Shi et al., 2022) have shown that the comments in source code are very noisy, with only a small fraction of comments actually describing the functionality of the code. The majority serve other purposes, such as licensing information, TODOs, or internal team coordination, making them poor indicators of intent. Relying on them alone leaves the model under-conditioned for real-world scenarios where intent is expressed ad-hoc. To close the gap, we need to supplement the model’s strong FIM foundation with targeted instruction alignment that empowers it to interpret explicit natural-language instructions.
@@ 2.3. Conversational LLMs are Sub-Optimal Code Completers
Conversational LLMs, such as GPT-4o, Claude-3, and DeepSeek-R1, have achieved remarkable performance on code-related dialog tasks. These include answering debugging questions, generating entire programs from natural-language specifications, and refactoring code snippets. In these scenarios, the model receives a textual instruction and produces a helpful, conversational response that mixes text and code. This capability is cultivated through an instruction tuning stage, where models learn to follow human instructions and dialog formats, often through supervised learning on instruction-response pairs (Wei et al., 2021). However, the datasets used for this purpose (Wang et al., 2024; Wei et al., 2023; Luo et al., 2023) primarily emphasize generating complete, assistant-style outputs rather than context-aware code insertions. This focus comes at a cost: the training process largely overlooks the strict contextual and logical consistency required by FIM code completion, where new code must integrate seamlessly with its surrounding prefix and suffix. As a result, when applied to FIM tasks, these powerful conversational models exhibit various failure modes, such as repeating the suffix or generating irrelevant content (Ahmad et al., 2025; Gong et al., 2024; Pan et al., 2024). An FIM-trained base model like CodeGemma-7b (Team et al., 2024) can easily outperform the much larger conversational LLM, GPT-4o, on real-world code completion benchmarks (Pan et al., 2024). Consequently, the LLMs deployed in production code completion systems are typically pre-trained base models, rather than their instruction-tuned counterparts.
@@ 3.3. Usage for Code Completion
An IFIM-trained LLM enables a developer to guide code completion by embedding instructions directly within their code. Notably, IFIM does not require developers typing instructions for every single completion they want. Instead, it serves as an effective solution to mitigate the unhelpful code completions caused by the under-specified developer intent. Since plain text would violate the programming language’s syntax, we propose writing instructions as specially marked inline comments. An example of this is shown in Figure 3. In this example, a developer using Python could write “#!filter out negative numbers” as the instruction. The ‘#!’ prefix acts as a unique marker that distinguishes an IFIM instruction from a regular comment. This allows client-side tools, like IDE plugins, to build user-friendly features around it. For example, an IDE could be configured to automatically remove the instructional comment once the generated code is accepted by the user. When a code completion request is made, the code completion system parses the code around the cursor to extract the prefix and suffix, and identifies the specially marked comment as the instruction. These three components, prefix, suffix, and instruction, are then formatted into the required IFIM input sequence. The IFIM-trained LLM receives this input and generates the code to fill the middle, which is then presented to the developer as a completion suggestion.
@@ 4. Experiment Setup
  • RQ4: Is representing instructions as a dedicated component effective?
  • RQ5: To what extent does the code generated by IFIM-trained models align with completion instructions?
@@ 4.2. Code LLMs
We evaluate our approach using two popular code LLMs as base models: Deepseek-Coder (Guo et al., 2024) and Qwen-2.5-Coder (Hui et al., 2024). Each model family includes a base version (pretrained with FIM objectives) and an instruction-tuned version (further fine-tuned on proprietary instruction datasets). In our experiments, we apply IFIM training to the base versions since they are the LLMs used for code completion. For a comprehensive comparison, both the original base models and their instruction-tuned counterparts serve as baselines in our evaluation.
@@ 4.3. Evaluation Benchmarks
Notably, these two benchmarks are not natively paired with code completion instructions. To evaluate the instruction-following capability of the models, we synthesize instructions for each sample using OpenAI GPT-4.1 (gpt-4.1-2025-04-14) with the same prompt employed in our instruction synthesis process. To further ensure the quality of the instructions, we manually revise those of low quality. Given the large sample sizes, we randomly select statistically representative subsets from each benchmark (95% confidence level, 5% margin of error, 50% population proportion), resulting in 312 and 256 samples for HumanEval-infilling and RepoMasterEval, respectively. The manual filtering process involves two authors independently reviewing all generated instructions and revising those that inadequately describe the desired code functionality. The authors achieve an inter-rater agreement of 0.72 and collaboratively resolve disagreements through discussion. The final curated instruction benchmarks are respectively named as IHumanEval and IRepoMasterEval (IRME), which are used in the experiments. We also publicly release these two derived benchmarks to facilitate future research.
@@ 5.1. RQ1: Effectiveness of IFIM
In this RQ, we systematically evaluate the effectiveness of IFIM, where we aim to assess two aspects: (1) whether IFIM can improve the instruction-following ability of the LLMs, and (2) whether IFIM can maintain the code completion ability when no instruction is given. For each base LLM (Deepseek-coder and Qwen2.5-coder), we compare its performance against two variants: (1) their official instruction-tuned version (instruct) and (2) our IFIM-trained version (ifim). The performance of these models are evaluated with the two benchmarks, IHumanEval and IRME, under two settings: (1) inputs with instructions and (2) inputs without instructions. In the “with instructions” setting, we pair each completion request with its corresponding instruction. For the base models, instructions are written as a comment appended to the end of the prefix. For the IFIM variants, instructions are constructed following the IFIM format. For the instruct variants, we embedded the instruction within a custom prompt designed for the fill-in-the-middle task. [^1] The prompt is available in our artifacts. In the ”without instructions” setting, inputs for the base and IFIM models are prefixes and suffixes without modification. For the instruct variants, however, we use the same prompt that asks the model to perform the fill-in-the-middle task, but without providing the instruction. The Pass@1 score is computed to assess the quality of the completion.
@@
Pass@1 Pass@1
Inputs with instructions Deepseek-Coder base PMS 84.6% 10.90%
instruct prompt 39.4% 2.10%
ifim PIMS 93.6% 21.10%
Qwen2.5-Coder base PSM 91.0% 18.40%
instruct prompt 21.8% 3.60%
ifim PSIM 95.8% 20.30%
Inputs without instructions Deepseek-Coder base PMS 68.6% 7.40%
instruct prompt 29.8% 0.60%
ifim PIMS 78.2% 16.00%
Qwen2.5-Coder base PSM 76.0% 10.20%
instruct prompt 20.8% 1.90%
ifim PSIM 76.3% 13.30%
@@
Moreover, the performance of instruct models aligns with the results in previous studies (Ahmad et al., 2025; Gong et al., 2024; Pan et al., 2024), i.e., existing instruction tuning methods impair their capabilities of fill-in-the-middle code completion. This is evident in both settings. When instructions are provided via a prompt, the instruct models perform poorly, with scores dropping to as low as 39.4% (Deepseek-Coder) and 21.8% (Qwen2.5-Coder) on IHumanEval. The degradation is also severe in the infilling setting without instructions. For instance, Deepseek-Coder-instruct’s score drops from a base of 68.6% to 29.8% on IHumanEval and from 7.4% to a mere 0.6% on IRME.
@@ 5.2. RQ2: Impact of IFIM Mode
Notably, the I-before-M modes (i.e., placing the Instruction component immediately before the Middle component) consistently achieve the highest instruction-following performance compared with other IFIM modes. For Deepseek-Coder, the PIMS mode is the top performer on both IHumanEval (93.6%) and IRME (21.1%). Similarly, for Qwen2.5-Coder, the PSIM mode achieves the best scores on both IHumanEval (95%) and IRME (20.3%). This demonstrates that placing the instruction immediately before the code to be generated most effectively enhances the LLMs’ understanding and execution of the given instructions. We therefore adopt I-before-M as the default mode for IFIM training.
@@ 5.4. RQ4: Impact of Standalone Instruction Component
Compared to treating instructions as a separate, standalone component within the format, a simpler, more straightforward solution would be to train LLMs directly using instructions embedded as in-line comments. In this RQ, we seek to investigate the impact of training on this newly introduced Instruction component, compared with simply learning instruction through inline comments. Specifically, we derive two training datasets from our instruction dataset: one processed using the IFIM format and the other by appending the instruction as an inline comment to the prefix. Using these two derived datasets, we train each base model (Deepseek-Coder and Qwen2.5-Coder) and correspondingly obtain two variants. Finally, we have two base models, each with one IFIM variant and one experimental variant for inline comments (named as CFIM, short for Comment-trained FIM). Since the instruction is appended as a part of the prefix, the inference for CFIM variants follows the same mode as their base models. We evaluate the performance of these models on IHumanEval-infilling and RepoMasterEval. During evaluation, the IFIM models are provided with instructed inputs, and the base model and CFIM models are provided with commented inputs (i.e., instructions are appended to the prefix as an inline comment). We compute the Pass@1 scores to measure the correctness of the completion.
@@
The results presented in Table 4 reveal that simply adding instructions as inline comments is an ineffective and often detrimental strategy. For Deepseek-Coder, this approach results in degraded performance, with Pass@1 scores dropping from 84.6% to 79.0% on IHumanEval and from 10.9% to a mere 4.3% on IRME. The Qwen2.5-Coder CFIM variant showed a similar, though less pronounced, negative effect, with performance decreasing on IHumanEval and remaining stagnant on IRME. This confirms a hypothesis: LLMs can be confused when instructional text is mixed with code as comments, negatively impacting their primary code completion capabilities. The stark contrast between the significant gains from IFIM and the performance degradation from CFIM suggests that the structural separation of instructions from the code, as implemented in IFIM, is important for the model to properly disambiguate, understand, and benefit from the provided guidance.
@@
5.5. RQ5: Alignment between Instructions and Code Completions
The IHumanEval and IRME benchmarks feature natural language instructions that are consistent with the surrounding code context. It could be questionable that the superior Pass@1 achieved by IFIM may come from the enhanced context understanding capability instead of the enhanced instruction-following capability. To this end, this RQ pushes the investigation further by understanding whether IFIM-trained models prioritize the natural language instruction over the implicit logic of the surrounding code. Specifically, we experiment to see when the intentions in the instruction and the code context are conflict, which side will the model follow. To create this conflict, we employed GPT-4 to systematically rewrite the instructions for each problem in the IHumanEval dataset. The rewriting process was guided by a specific prompt [^2] This prompt is available in our artifacts. that instructed GPT-4 to craft a new instruction that was deliberately different from the original one but still runnable in the code context. The written instructions are thus conflicting with the correct code expected by the code context. Some examples of these conflicting instructions are available in Figure 4.
Using these newly created conflicting instructions for IHumanEval, we evaluate both the base and IFIM versions of DeepSeek-Coder. Our goal is to evaluate instruction alignment, i.e., whether the generated code align with the provided instruction. Therefore, in addition to Pass@1, we perform an human annotation to judge the alignment between the generated code and the completion instructions. Specifically, following a similar protocol to our benchmark creation, two authors independently classified each code completion based on whether the code implements the logic from the instruction. Notably, the annotators are required to only consider the alignment of the code’s logic with the instruction’s intent, irrespective of the code’s functional correctness. The evaluators achieved an inter-rater agreement of 0.69, and all disagreements were resolved through discussion to reach a final consensus.
Table 5. Evaluating Instruction Alignment on Conflicting Instructions
Model Version IHumanEval (Conflicting Instructions)
Pass@1 Alignment
Deepseek-Coder base 84.6% $\rightarrow$ 59.6% 54.4%
ifim 93.6% $\rightarrow$ 56.7% 80.9%
We report the pass@1 and the ratio of examples that are annotated as aligned in Table 5. The most direct measure of this is the alignment score. The IFIM-trained Deepseek-Coder achieves an alignment of 80.9%, indicating that in a vast majority of cases, it correctly generated code that implemented the logic of the new conflicting instructions. In contrast, the base model only achieved 54.4% alignment. This suggests that the base model was often confused by the conflict, frequently ignoring the commented instruction and instead generating code based on the original function signature and context. This conclusion is further reinforced by the Pass@1 scores. On the surface, both models show a significant drop in performance against the original test cases. This is an expected outcome, as correctly following a conflicting instruction should logically lead to a failure on the original tests However, the drop is more pronounced for the IFIM model (from 93.6% to 56.7%) than for the base model (from 84.6% to 59.6%). This steeper decline for the IFIM model is a direct consequence of its superior instruction-following ability. By successfully implementing the new instructions more often (as confirmed by the 80.9% alignment score), it naturally failed the now-irrelevant original tests at a higher rate.
We further qualitatively demonstrate the instruction-following behaviors between base and IFIM models using five real examples, shown in Figure 4 In each case, the base model tends to ignore the commented instruction, defaulting to the logic suggested by the original function context. For instance, in Example 5 (HumanEval/63/L4), given the conflicting instruction that requires to check if n equals to 3, the base model still generates ‘if n == 2:‘, the code expected by the code context. Conversely, the IFIM model adheres to the new instruction, producing ‘if n == 2:‘. These examples provide qualitative evidence that IFIM training genuinely enhances the model’s ability to prioritize and follow explicit natural language instructions, even when they contradict the surrounding code context.
Figure 4. A visual comparison of outputs produced by base and IFIM DeepSeek-Coder given conflicting instructions. The checkmark () indicates that the generated code aligns with the intent of the Conflicting Instruction, while the cross mark () indicates that it does not.
Example 1: HumanEval/143/L3 Original Instruction: Check if the length of the word is 1.
Conflicting Instruction: Check if the word starts with ’a’.
Base model: if word[0] != ’a’: IFIM model: if word[0] == ’a’:
Example 2: HumanEval/140/L12 Original Instruction: Append the current non-space character.
Conflicting Instruction: Append an underscore character.
Base model: new_text += text[i] IFIM model: new_text += "_"
Example 3: HumanEval/3/L4 Original Instruction: Check if the balance is less than zero.
Conflicting Instruction: Check if the balance is greater than zero.
Base model: if balance < 0: IFIM model: if balance > 0:
Example 4: HumanEval/13/L0 Original Instruction: Repeat the loop as long as b is not zero.
Conflicting Instruction: Repeat the loop as long as a is not zero.
Base model: while b != 0: IFIM model: while a != 0:
Example 5: HumanEval/63/L4 Original Instruction: Check if n equals to 2.
Conflicting Instruction: Check if n equals to 3.
Base model: if n == 2: IFIM model: if n == 3:
Answer to RQ5: IFIM-trained models demonstrate strong instruction alignment, prioritizing explicit instructions over the intention form code context. The IFIM variant followed conflicting instructions in 80.9% of cases, a significant increase from the base model’s 54.4%, confirming that the performance gains are attributable to enhanced instruction-following capability.